OpenStudy (anonymous):

Please Help... Differentiate... Y= Srqt(1+ 2e^3x)

OpenStudy (anonymous):

\[Y= \sqrt{1+2e^3x}\]

OpenStudy (anonymous):

chain rule plus product rule \[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\] with \[f(x)=1+2e^{3x}, f'(x)=6e^{3x}\]